OpenStudy (calculusxy):

MEDAL!!! It takes me about an hour to get from my home to the supermarket. If my home is x km away, what is my average speed in meters per second? A. 1000x B. x/1000 C. x/1000 X 60 D. 1000x/60 E. x/1000 X 60 X 60 F. 1000x/60 X 60

OpenStudy (calculusxy):

@Nnesha

OpenStudy (astrophysics):

\[s = \frac{ d }{ t }\] s is your speed, d is your distance (x in this case) and t is your time.

OpenStudy (calculusxy):

I chose D, but I know I am wrong. Can you explain to me why?

OpenStudy (astrophysics):

Well your answer wants it in meters/ second so notice we are given 1 hour, and x km, so first thing we want to do is convert it to seconds and meters respectively. To convert an hour to seconds we do 60*60 which will give us our time in seconds. For meters, know that 1km = 1000m, so we have 1000x using our formula above, for speed we have \[s = \frac{ d }{ t } = \frac{ 1000x }{ 60 \times 60 }\]

OpenStudy (calculusxy):

My answer is correct? Are you sure because my teacher marked that wrong?

OpenStudy (astrophysics):

Sorry I thought you picked F

OpenStudy (astrophysics):

the last option

OpenStudy (astrophysics):

D would be wrong because, you have it in minutes not seconds

OpenStudy (calculusxy):

Oh yes, I didn't look for that. Thank you!

OpenStudy (astrophysics):

Np :)

OpenStudy (calculusxy):

I have another question.

OpenStudy (calculusxy):

I did an experiment today relating to torques. The materials that I had were a meter stick, masking tape, and weights. We had to use the ruler and place the two weights on each side, to try to make it balance to each other. Also, we had to note how far the tape (holding the weight) was from the center (or 50cm). Here is my first data: Mass 1: 200g Distance from 50cm: 19cm Mass 2: 100g Distance from 50cm: 36 cm Then we had to multiply the mass times the distance for each and were asked to look for a pattern. Here are my products: Mass 1: 200g x 19 cm = 3800 Mass 2: 100g x 36 cm = 3600 I don't notice a pattern and need help in finding one.

OpenStudy (calculusxy):

@Astrophysics